OpenStudy (anonymous):

i need help with n+5(n-1)=7

OpenStudy (saifoo.khan):

n + 5n - 5 = 7

OpenStudy (saifoo.khan):

6n = 7 + 5 6n = 12 n = 12/6 n= 2